Five years ago, John was three times as old as Lisa and in ten years from now John shall be two times as old as Lisa. What are their current ages?
Answer by htmentor(1343)

Five years ago, John was three times as old as Lisa and in ten years from now John shall be two times as old as Lisa. What are their current ages?
====================
Let J = John's current age
Let L = Lisa's current age
5 years ago John's age (J-5) was equal to 3 times Lisa's age (L-5), so
J-5 = 3(L-5) -> J = 3L - 10
In 10 years John will be twice as old as Lisa:
J+10 = 2(L+10)
Using the expression above for J in terms of L, we have
(3L-10) + 10 = 2(L+10)
3L = 2L + 20
L = 20
So Lisa is 20 yrs old and John is 3*20 - 10 = 50 yrs old